# Enhancing Photon Identification with Neural Network Methods

### v1

@@ 5 Conclusions
In this study, we generated and simulated photon and pion samples using the COCOA-HEP tool, a state-of-the-art detector simulation framework. We emulated the current baseline method for photon identification, a BDT classifier based on shower-shape variables, and introduced DNN and ResNet architectures as the natural next step toward improved performance.
Our results demonstrate that the ResNet significantly outperforms both the BDT and DNN across key metrics, particularly in the very low false-positive rate regime relevant for high-purity photon identification. Furthermore, we showed that fine-tuning with soft scoring and incorporating angular separation information leads to additional gains, enabling the model to better discriminate the most challenging background cases without sacrificing signal efficiency.
These findings highlight the potential of modern deep learning architectures, combined with physics-informed loss functions, to improve photon identification in collider experiments. Future work could extend these methods to full detector simulations and explore applications for real experimental data.
In contrast to previous deep-learning-based photon identification studies, which largely emphasized global shower properties or jet discrimination, our approach explicitly targets the substructure of overlapping electromagnetic showers in a symmetric calorimeter geometry. This method provides a clean, controlled benchmark for future photon identification developments at the High-Luminosity LHC and beyond.

### v2

@@ 5 Conclusions
In this study, we generated and simulated photon and pion samples using the COCOA-HEP tool, a state-of-the-art detector simulation framework. We emulated the current baseline method for photon identification, a BDT classifier based on shower-shape variables, and introduced DNN and ResNet architectures as the natural next step toward improved performance. Our results demonstrate that the ResNet significantly outperforms both the BDT and DNN across key metrics, particularly in the very low false-positive rate regime relevant for high-purity photon identification. For example, ResNet achieves $E_{90}=79.3$ GeV while neither baseline method reaches 90% efficiency within the studied $p_{T}$ range. We then systematically explored two approaches to push performance further: soft scoring (assigning continuous labels based on $\Delta R$) and an auxiliary $\Delta R$ regression head. Soft scoring with a linear function peaked at $E_{90}=74.0$ GeV, while the auxiliary head with $\alpha=0.75$ achieved $E_{90}=74.5$ GeV—both representing substantial 6% gains over the ResNet baseline.
Interestingly, combining the methods proved counterproductive, yielding $E_{90}=75.3$ GeV. This dip in performance relative to the individual models indicates a level of mutual interference, where the two strategies compete for the same network resources rather than complementing one another, and might require a simultaneous hyper-parameter scan for optimization. We also tested robustness under systematic uncertainties. Energy scale miscalibrations up to ±5% caused minimal degradation ( 1 GeV), but excess calorimeter noise proved more challenging. Adding 20 MeV Gaussian noise degraded the baseline by over 10 GeV, while soft scoring held up remarkably well with only 4.5 GeV degradation. The auxiliary head showed no such advantage, degrading by 11.3 GeV. This makes soft scoring particularly attractive for real detector conditions where noise levels fluctuate.
These findings highlight the potential of modern deep learning architectures, combined with physics-motivated training strategies, to improve photon identification in collider experiments. Future work could extend these methods to full detector simulations and explore applications for real experimental data. In contrast to previous deep-learning-based photon identification studies, which largely emphasized global shower properties or jet discrimination, our approach explicitly targets the substructure of overlapping electromagnetic showers in a symmetric calorimeter geometry. This method provides a clean, controlled benchmark for future photon identification developments at the High-Luminosity LHC and beyond.